Large Binocular Telescope Commissioning - Focal Stations

Background information, manuals, etc. on:

LBTtoolsIRTC (IRAF Observing Interface) and JohnSpecial (IRAF and Python scripts)
RunningTheIRTC (IRTC Engineering Interface) and RunningDirtI (IRC/IRS Socket Interface)
Running the AGwStages (AGw Engineering Interface), Running the AzCam Server, Status for AGwUnits
RunningGCS (Guiding Control), RunningPCS (Pointing Control), BinocularObserving (Mixed Instruments)
NonSiderealGuiding, RunningPSF (Point-Spread Function), and RunningOSS (Optical Support)
ByEyeAO (active optics by inspection), ActiveOpticsModes (PSF arbitrator control)
RunningEF (Analysis of Intra-/Extra-focal Pupil Pairs)
CollimationModel and PointingModel
AGW Coordinate Transformations (481g221c, updated 10-Oct-2007)
VNCforDummies (instructions to mirror the workstation display downtown)
BeginningOfTheNightIRTC (startup cookbook) and BGTechProcedures (procedures for AGw alignment and such)
AOwithIDL (active optics with the AGw S-H sensor)
Restarting the IDL license Manager
RotatorAngle (notes about rotators and other interesting angles)
Adaptive Secondary seeing limited instructions

User Information

BeginningOfTheNightLUCIFER (startup cookbook)
Focal Station Commissioning Tools (IDL) (procedures for various commissioning measurements)
IDL Commands
LBTO Star Catalogs
WhereIsIt (where are data and log files located) ImportantParameters (for LBTtools with LFBG/RFBG)
GenericLuciferTechNightsPlan
WhatIsTheSeeing
MODS off- and on-sky checkout MODS tools: slitscan (obsolete) and modsastrom (added 2015-08-04 by Olga Kuhn)

Request for On-Sky Use of the Telescope

Form for submitting Eng On-Sky Requests

Form for submitting D-time On-Sky Requests (LBTO email required)

Technical Observing Logs

Note: These logs are mainly for telescope, eng/D-time. Additional logs may be found under Instruments and Adaptive Optics sections of the LBT wiki.

2023BTechnical Observing Logs

29 Nov 2023 - ECD: PEPSI POL checkout, other

25 Nov 2023 - ECD 1/2: Split LBTI/ECD (Weathered out)

24 Nov 2023 - ECD: SOUL Comm and DDT

01 Nov 2023 - ECD: LUCI1 Checkout and LUCI2 AO

31 Oct 2023 - ECD: LUCI1 Checkout and MODS transform

18 Oct 2023 - ECD: LUCI2 Checkout

05 Oct 2023 - ECD: Chiller maintenance

23 Sept 2023 - ECD: LBTI and AO restart

22 Sept 2023 - ECD: LBTI and AO restart

10 Sept 2023 - Restart Night9:

09 Sept 2023 - Restart Night8:

08 Sept 2023 - Restart Night7:

07 Sept 2023 - Restart Night6:

06 Sept 2023 - Restart Night5:

05 Sept 2023 - Restart Night4:

04 Sept 2023 - Restart Night3:

03 Sept 2023 - Restart Night2:

02 Sept 2023 - Restart Night1:

2023 Restart Plan - Planning Page for Restart

2023A Technical Observing Logs

29 June 2023 - ECD - MODS DDT & phot imaging field; Shark NIR offset testing; LBC phot standards, short time on Z22 and thermal pupil data collection; seeing variable and up to about 2".

26 June 2023 - ECD Split w/INAF: tested TMS/FPIA automation; LBC phot stds; checked LUCI2 N30 focus DIQ @ FeII simul w PSF morphology

05 June 2023 - ECD: SN with PEPSI, PSF reconstruction, verification of field aberration.

27 May 2023 - ECD: Field aberration data collected, analyzed, implemented. DDT programs using PEPSI & MODS; MODS Spectrophot (for M1:M2 comparison, not photometric)

26 May 2023 - First night after the rigid M2 was installed on SX: MODS Spectrophot & Imaging standards; DDT PEPSI & MODS; ESM/SESM

24 May 2023 - First Half ECD with PEPSI and SHARKNIR

04 Apr 2023 - Split night w/LBTI

03 Apr 2023 - DEC Night - D-time observation, LUCI-AO task, SOUL-LUCI2 commissioning

21 Mar 2023 - DEC Night - The entire night was lost due to bad weather.

06 Mar 2023 - DEC Night - AGw2 focus check on guider (prelim results suggest current value ok, additional test tbd); LBTI optimize gain checks

27 Feb 2023 - DEC Night - Snow and ice on the enclosure prevented opening.

10 Feb 2023 - DEC Night - MODS+LUCI (spectro-)photometric standards, LBC TMS on-sky channel dropping tests, DDT observation AGw patrol field check

04 Feb 2023 - DEC Night - DX adSec back. ESM bin 1 binocular testing (first DX ESM). LBC TMS on-sky channel dropping tests, though temperature swings caused problems for FPIA collimation and the tests were limited.

03 Feb 2023 - DEC Night - DX adSec not available. ESM testing (gain ranges/routine tuned) on SX. SL on DX. Bins 1 & 4. Investigated AGw1 patrol field vignetting (IT 8495). Updated FLAO (not really FLAO, historic) TTM-Auto acquisition routine to include optimize gain.

02 Feb 2023 - DEC Night - DX adSec taken offline ~10UT (IT 8825). SX-SOUL optimize gain testing. TCS and AO software updates deployed and, seeing no problems, left in place.

2022B Technical Observing Logs

13 Jan 2023 - Daytime - AGW2 guide probe to rotator center

01 Jan 2023 - Split Night - LBTI 1st half / LBTO 2nd half

31 Dec 2022 - DEC Night (1st half LBTI / 2nd half LBTO) - no chance to open due to weather - closed dome SuperESM tests run 2nd half

25 Dec 2022 - DEC Night - shaky seeing at start - Most D-time completed, several SuperESM tasks run

11 Dec 2022 - Split Night w/PETS - MODS restart completed

09 Dec 2022 - Split Night - LBTB/LN-Com first half, Eng 2nd half (MODS restart) - Night lost to weather, closed-dome TMS testing & MODS cals.

08 Dec 2022 - DEC Night - Night lost to weather, no chance to open

07 Dec 2022 - Split night w/PEPSI-Pol - Night lost to weather, no chance to open

06 Dec 2022 - Split night w/PEPSI-Pol - Night lost to weather, no chance to open

02 Dec 2022 - Split Night w/PETS-PEPSI - some time lost to clouds and winds - TMS testing done, channel dropping work still needed, no chance to execute D-time science

17 Nov 2022 - DEC night - late opening for PEPSI-Pol issue, poor/variable seeing plagued most of night, PEPSI-Pol checkout/calibrations completed, some AO/ESM work attempted

05 Nov 2022 - Split Night w/LBTI Science - some variable seeing during night, otherwise successful AO eng/D-Time programs executed

04 Nov 2022 - Split Night w/LBTI Science - DEC 2nd half - Night called for ice before sunset

03 Nov 2022 - Split Night w/LBTI Science - DEC 2nd half - Night lost to weather

16 Oct 2022 - Split Night w/PETS - Night lost to weather

11 Oct 2022 - Split Night w/LBTI (1st half), DEC (2nd half) - No chance to open, night lost to weather

09 Oct 2022 - Split Night w/LBTI (1st half), DEC (2nd half) - No chance to open, night lost to weather and internet outage

08 Oct 2022 - Split night w/PETS - No chance to open, night lost to weather

04 Oct 2022 - Split night w/PETS - No chance to open, night lost to weather

29 Sep 2022 - Split night w/PETS - installed AGW1/2 transforms permanently, collected field aberration confirmatory data (and processed), AO checks - first attempt failed, later did Bin 3 and OSCO, Bin 4 failed. No ESM tests.

25 Sep 2022 - Night #14 of Restart - split night w/PETS - MODS critical checkout mostly completed, AGW1/2 transform confirmatory data collected

22 Sep 2022 - Night #13 of Restart - rollover of MODS checkout on-sky, AGW1/2 Confirmatory Transform, Field Aberration Confirmatory data collection - Night lost to weather - telescope at horizon all night - very limited MODS checks done

21 Sep 2022 - Night #12 of Restart - MODS checkout on-sky, AGW1/2 Confirmatory Transform, Field Aberration Confirmatory data collection - Night lost to bad weather, critical issues with MODS, no MODS checkout or calibs

13 Sep 2022 - Night #11 of Restart - split night w/PETS - Field Aberration Data, AGW Transforms, LBC/TMS priorities - no chance to open, Night lost to bad weather

11 Sep 2022 - Night #10 of Restart - Rollover of previous tasks - Field Aberration Confirmatory data, LUCI checkouts, TMS/LBC Checkout - most of night lost to weather - field aberration unsuccessful, G210 LUCIs tested, troubleshooting TMS

10 Sep 2022 - Night #9 of Restart - Rollover of previous night tasks - LUCI checks, AGW1/2 checks, AO if time, LBC TMS testing (if system is ready) - Night Lost - no chance to open - AO/ESM closed dome checks run

09 Sep 2022 - Night #8 of Restart - Rollover of previous night tasks - LUCI checks, AGW1/2 checks, AO if time - Night Lost - no chance to open due to thick clouds - LUCI flexcomp cals run

08 Sep 2022 - Night #7 of Restart - split with PETS - Night Lost - No chance to open due to thick clouds - LUCI flexcomp cals run

07 Sep 2022 - Night #6 of Restart - Continue with LUCI checkout - LBC skyflats in evening, RotZP checks AGW1/2, AGW 1/2 Transform collect, AO on-sk, major problem with LUCI sw (unusable for science),

06 Sep 2022 - Night #5 of Restart - Rollover of tasks, Field Aberration Data (AGW1/2), LUCI checkout - completed collecting Field Aberration data, collected LUCI rotator data and RotZP for LUCI-2 w/MOS, tested LUCI NS, then seeing blew up.

05 Sep 2022 - Night #4 of Restart - Rollover of tasks, AGW1/2 rotator hotspot, M3 alignment, Field aberration data, start of LUCI checks - Completed PEPSI-Pol checkouts, AGW1/2 RotHotspot, M3 Pupil Wobble (both sides), LBC checks in poor/variable conditions

04 Sep 2022 - Night #3 of Restart - Rollover of any tasks not completed, PEPSI Flux Cals (Pol/PFU), LUCI wakup - PEPSI-POL SX checkout, rotator hotspot AGW1, lost 6+hours to weather

03 Sep 2022 - Night #2 of Restart - AGW1/AGW2 checkout, Telescope Wakeup continued, PEPSI Pol/PFU Checkout - slew of critical ITs, PEPSI-PFU released for science, lost several hours to bad weather

02 Sep 2022 - Night #1 of Restart - Telescope wakeup, LBC checkout, AGW1/2 checks - LBC checkout partially completed - problem with U-Bessell filter on LBC-Blue - lost several hours due to clouds and high humidity.

2022 Restart Plan - Planning Page for Restart

2022A Technical Observing Logs

14 Jul 2022 UT - - Extra Engineering Night - SX ASM removed, replaced with rigid secondary - compare flux with DX ASM (redux)

13 Jul 2022 UT - Extra Engineering Night - SX ASM removed, replaced with rigid secondary - compare flux with DX ASM - No chance to open for scary clouds be above us all night!

08 Jun 2022 UT - Full DEC night - ESM tests, Bin 1 Full AO testing (GOpt, optimize gain, NCPA) to narrow down triggers of problems, seeing stable all night, but clouds thickened last few hours (still good seeing)

07 Jun 2022 UT - Full DEC night - MODS science both completed, OSCO failure, setz0 test worked for Bin2 and Bin3, attempted Bin1 at end of night, to be continued next night

29 May 2022 UT - split night with PETS/PEPSI - DEC until 21:35, PEPSI/PETS and PEPSI Eng after until sunrise - DEC time lost to winds - PETS time lost to winds - PEPSI Eng flux cal standard completed

14 May 2022 UT - DEC Night - LBTI until 1am (local time), ESM characterization after (faint end WFSMag to 18.3, IQ testing, bright bin 1 testing)

13 May 2022 UT - split night with LBTB (1st half), DEC (2nd half) - LBTB science w/ESM SUPER successful, LBTI used 2nd half for COM/Eng

09 May 2022 UT - DEC Night -lost most of the night to high winds - some attempts at AO, bayside stage test completed

08 May 2022 UT - Split Night with PETS/PEPSI - PETS 07:55-10:25 (UT - including reconfig time) - LBTI DX work successful, lost time for PETS to winds, fell back to poor seeing programs for rest of DEC time

05 May 2022 UT - Split Night with PETS/PEPSI - PETS 09-11:40 (UT) - poor seeing - some LBC/TMS testing done

26 Apr 2022 UT - Split Night with PETS/PEPSI - PETS twilight-09:21 (UT) - extremely variable seeing all night - attempted (failed) OSCO, offsetxy AO test successful, fell back to backup science end of night

18 Apr 2002 UT - DEC Night - TMS/LBC testing, some D-time w/LBC, AO/ESM tests (faint WFS mag stars, Bin2 ESM Bin2 Full AO)

17 Apr 2022 UT - DEC Night - highly variable seeing all night - OSCO done, some ESM testing, fell back to LUCI photometric standards and poor conditions science program

16 Apr 2022 UT - Split Night - LBTB Partner Science / DEC - thick clouds (w/gusts) most of night - break in clouds allowed field aberration data to be collected, 1 poor weather object attempted at end

08 Apr 2022 UT - DEC Night - LBC/TMS testing, short AO use, some D-time - highly variable seeing all night

07 Apr 2022 UT - Split Night with PETS/PEPSI - DEC 1st part, PETS/PEPSI 2nd half

02 Apr 2022 UT - Split Night with PETS/PEPSI - DEC 2nd part - D-time tasks mostly completed (did not complete one pass of the LBC observations), MODS D-time science completed.

14 Mar 2022 UT - Full DEC Night (iLocator may use 0.5night if cannot use previous two nights) - TMS testing first half, poor seeing all night, LBC phot standards, LUCI phot standards, then closed early for winds

13 Mar 2022 UT - Possible Dec time (if conditions too poor for iLocator) - continued iLocator testing (successful), handed over to AO - Bin3 tests, repoint tests, binary separation test, then back to AZ queue

12 Mar 2022 UT - Possible Dec time (if conditions too poor for iLocator) - rollover for 13 Mar 2022 UT - iLocator testing successful - handed back to AZ science 1hour early to take advantage of excellent conditions - no DEC executed

06 Mar 2022 UT - Split Night with PETS/PEPSI - DEC time - no chance to open

05 Mar 2022 UT - Split night with PETS/PEPSI -DEC time - night called for rime ice - closed dome AO testing

20 Feb 2022 UT - Split Night with PETS/PESI - DEC time - useful AO testing

18 Feb 2022 UT - Split night LBTI 1st half - DEC 2nd half - some AO eng tasks attempted, highly variable seeing

16 Feb 2022 UT - Split night LBTI 1st half - DEC 2nd half - Closed for winds almost all night - two brief windows to open, couldn't complete collimation before multiple 22m/s gusts hit - tasks rolled over for Eng

15 Feb 2022 UT - Split night LBTI 1st half - DEC 2nd half - Eng tasks completed

14 Feb 2022 UT - Full DEC night - 1st half LBC/TMS - 2nd half LUCI-AO - very successful 1st half, 2nd half some issues with Bin 1 ESM (filter issue on OCAM and re-rotator limit), Bin 2 ESM, Bin 3 Full AO IQ - characterized "jumps" on LUCI images.

08 Feb 2022 UT -split night PETS/DEC - LUCI-AO testing 2nd half - collected Bin 1 and BIn 3 data for several stars - seeing started at 5" then dropped to less than 1"

07 Feb 2022 UT - split night PETS/DEC - LBC TMS testing (extrafocal tests in poor seeing), no chance for LBTI, post-PETS seeing 5"+ could not maintain any guiding, let alone WFS'ing

04 Feb 2022 UT - Split Night PETS/DEC - No chance to open - UPS/Electrical issues put instrument/telescope at risk

03 Feb 2022 UT - Split Night PETS/DEC - No chance to open - UPS/Electrical issues put instrument/telescope at risk (poor weather as well)

02 Feb 2022 UT - Split night PETS/DEC - No chance to open - UPS/Electrical issues put instrument/telescope at risk (poor weather as well)

Feb D_Eng_Com Plan - Landing page for UT Feb DEC nights

2021B Technical Observing Logs

22 Jan 2022 UT - D/E/C night - MODS checkout completed (hotspot w/rt rotator center, RotZP, SpecStd all modes), LUCI-1 AO/ESM (Bin 1,2,3,4 / ESM Bin 4)

21 Jan 2022 UT - D/E/C night - LBC/TMS testing, 1994PC1 flyby observations (poor seeing), LUCI-1 AO on-sky checkouts

15 Jan 2022 UT - Split Night w/LBTI - DEC 2nd half - poor conditions - some PEPSI/Pol work done

29 Dec 2021 UT - Split Night PETS-PEPSI/DEC - twlight-02:25hr PETS, DEC after - no chance to open - night called for snow, ice, high winds, etc.

27 Dec 2021 UT - Split Night w/PETS-PEPSI - 1st haf DEC - ADC work done (more needed), runaway Zernikes corrected, closed before PETS science, never reopened

23 Dec 2021 UT - Split Night with INAF - DEC 2nd half (PEPSIPOL checkout, AGW3/4 checkout, clouds) - AGW4 checks completed, some work on zernike and ADC

16 Dec 2021 UT - Split Night Eng/PEPSI/LBTI - No chance to Open - dome encased in ice.

13 Dec 2021 UT - DEC Time split with LBTI - 1st half LBTI-Eng and LUCI-1 AO Eng - 2nd half LBTI science - TMS tested with LBCs successfully

02 Dec 2021 UT - DEC time - 3hrs - end of night - ESM bin 1 OK, ESM bin 4 did not work, ESM bin 2 did not work - LUCI2- RotZP and transform check run

01 Dec 2021 UT - AO Eng 2.5 hours start of night to verify ESM - AO tests ran, ESM test did not work (AO Ref star too faint)

28 Nov 2021 UT - Split Night - D/E shared w/PEPSI/PETS - PETS UT 03:30-08:36 UT (08:30-01:36 AZ time) - open all night, DEC time for LUCI-1 AO successful - some issue remained for dayytime work

24 Nov 2021 UT - D/E night - no chance to open - closed dome SX LUCI-AO work

23 Nov 2021 UT - D/E night - opened a short time - completed 1 OSCO (no-NCPA), closed dome testing of AO Bin 1 and Bin 4 after.

20 Nov 2021 UT - Split night - D/E shared w/PEPSI/PETS - no AO executed due to techincal issues, backup D, PET/PEPSI, GRB followup D-time

19 Nov 2021 UT - Split night with LBTI (2nd half), D/E (1st half) -M31 Nova data, LBC Timing tests

18 Nov 2021 UT - D/E night - successful Pallas observations with LBTI, M31 nova, LUCI sub-frame readout testing, LUCI science (1hr), LBC timing tests w/TMS

17 Nov 2021 UT - D/E night - Pallas attempt w/LBTI, M31 nova, IT 8472 MODS checks, TMS 2nd half

16 Nov 2021 UT - D/E night split with PEPSI -M31 nova, LUCI sub-frame testing, PEPSI, merger MODS - successful night

15 Nov 2021 UT - D/E night - LBTI until 20:30, TMS/LBC until 00hr, Science after - LBTI eng completed, science tasks completed, TMS eng progress made (additional time will be needed)

Nov_Eng_D_Plan - Landing page for UT Nov 15-20, 23-24

09 Nov 2021 UT - Split with PEPSI/PETS - D/E from twilight until UT 04hr (21hr local time) - AGW2 Transform confirmatory data collected, LUCI PhtStd

04 Nov 2021 UT - Split with PEPSI/PETS - D/E from 06:30 UT (23:30 local time) onwards - successful LBTI tests on-sky, successful TMS testing in dark/good conditions

31 Oct 2021 UT - split night PEPSI/PETS until 09:10UT - D/E after - LUCI-2 AGW transform obtained

27 Oct 2021 UT - D/E Night - poor seeing but clear - completed MODS standards, photSpec standards in all gratings/prism combinations, MODS WFS cutout issue tested, TMS testing with LBCs

26 Oct 2021 UT - D/E Night - opened brielfy - some LBTI AO testing done - no LBTI science. Closed rest of night, no other tasks started or completed.

01 Oct 2021 UT - Split Night - PEPSI/PETS until 07:00 UT - D/E after

29 Sep 2021 UT - Split Night - PEPSI/PETS until 07:44 UT, D/E after - No chance to open

16 Sep 2021 UT - Restart Night #13 - MODS science observations, MODS skyflats, Spec Std, dither test, LUCI-1 Rot ZP verificaiton, LBC skyflats

15 Sep 2021 UT - Restart Night #12 - Poor seeing most of night (as high as 4") - PEPSI training, LBC Pointing Model (w/active TMS), LBC Std fields (w/active TMS), LBC skyflats

14 Sep 2021 UT - Restart Night #11 - LUCI Stress Test, Confirmatory AGW-1 Transform, NS dither test, LBC Pointing Model

13 Sep 2021 UT - Restart Night #10 - R.Oph monitoring, confirmatory MODS AGW Transform, check on Rot ZP and dither, Skyflats LBC - issue with bug goo on MODS-2, WFS cutouts on both MODS offset, lost time for smoke

12 Sep 2021 UT - Restart Night #9 - first few hours some science (since previous night was used entirely for Eng), Field Aberration confirmatory data (via LFBG, RFBG), NS LUCI-1 LS check, Mixed mode (MODS-1 & LBC-R)

11 Sep 2021 UT - Restart Night #8 - PEPSI/PETS until UT 06:43 (11:43 local); D/E-Restart after - PEPSI pointing model, M31 LUCI-1 offset test, LUCI skyflat

08 Sep 2021 UT - Restart Night #7 - MODS checkout continued (longslit acq, blind offset, NS imaging); LUCI-1 checks continued (G210, G200 longslit)

07 Sep 2021 UT - Restart Night #6 - LBC Skyflats (PA=180, 6 filtes), MODS on-sky checks (RotZP, Field Center, AGW Transform, Dither test) - LBC-R 0 count error, MODS2 mask error

06 Sep 2021 UT - Restart Night #5 - Completed Field Aberration collection for SX and DX; LUCI-1 tasks for: field center, Rotator ZP, AGW Transform, Imaging tests (Sidereal, NS), Photometric Std.

05 Sep 2021 UT - Restart Night #4 - opened 1hr end of night - PEPSU PFU AGW checkout, attemped field aberration collection (data probably not usable)

04 Sep 2021 UT - Restart Night #3 - No chance to open - completed MODS2 Grating and Imaging Cals

03 Sep 2021 UT - Restart Night #2 - AGW 1&2, AGW 5&6 alignment and checks, M3 alignment data collected, Field aberration collected (except DX w/AGW2), Main LBC restart tasks executed

02 Sep 2021 UT - Restart Night #1 - closed most of the night - 1.5hrs of LBC initial checks

2021 Restart Plan - Planning Page for Aug/Sep restart

2021A Technical Observing Logs

12 Jul 2021 UT - Split night with PETS - handover to PETS @21:25 local time

28 Jun 2021 UT - Split night with PETS - handover to PETS @22hr local time

25 Jun 2021 UT - Engineering Night for AO - 1st half LBTI, 2nd half LUCI

11 Jun 2021 UT - D/E half night - 1st, 2nd half PEPSI

03 Jun 2021 UT - D/E half night - 1st half

31 May 2021 UT - D/E half night 1st half (2nd half LBTI Science)

29 May 2021 UT - MODS Checkout (up to 0.5 night)

24 May 2021 UT - AGW4 checkout at the start of PEPSIPOL

22 May 2021 UT - AGW4(and 3) PEPSI-Pol checkout (lost to weather)

21 May 2021 UT - D/E night - 1st half LBTI-SOUL-Com, 2nd half Eng (primarily PEPSI-Pol checkout)

May_Jun_2021_ED_Plan - Landing Page for May, Jun, Jul D/E time

01 May 2021 UT - few hours LUCI-ESM Eng

29 Apr 2021 UT - Engineering Night (Full)

29 Mar 2021 UT - D/E Full night

14 Mar 2021 UT - split night with PEPSI/PETS - MODS-Bino before, LUCI-Bino/ESM after - No chance to open - snow storm!

05 Mar 2021 UT - split night with LBTI - D/E 2nd half - on-sky data/science collected

04 Mar 2021 UT - split night with PEPSI/PETS - D/E before and after - D/E portions weathered out

03 Mar 2021 UT - split night with PEPSI/PETS (1st half) - 2nd half AO-SOUL-LUCI-1 on-sky testing - AO testing collected good quality data

02 Mar 2021 UT - split night with LBTI - LBTI first half - D/E - AO SOUL-LUCI-1 on-sky testing 2nd half - no LBTI due to UBC failure, poor seeing (2-5.6" on guider) almost all night - 1hour of AO valiantly attempted at end

01 Mar 2021 UT - split night with LBTI - LBTI first half - D/E - AO SOUL-LUCI-1 on-sky testing 2nd half - SX M3 could not move to LUCI-1

28 Feb 2021 UT - split night with LBTI - D-time first 3hrs w/LBTI, then LBTO-Eng (MODS), 2nd half LBTI Science - AO issues with LBTI, then winds closed before MODS Science/tasks.

27 Feb 2021 UT - split night with LBTI - D-time first 2hrs w/LBTI, then LBTO-Eng (LBCs TMS), 2nd half LBTI Science

26 Feb 2021 UT - split night with LBTI - D-time first 3hrs w/LBTI, then LBTO-Eng (MODS), 2nd half LBTI Science- 1st half hit by very poor seeing (2-5")

25 Feb 2021 UT - split E/D wth PEPSI (non-PETS) - handover at night center - E/D 2nd half - top priority LUCI-1 AO Engineering - PEPSI lost time to winds, AO/Eng time impacted by variable seeing, clouds and Moon

14 Feb 2021 UT - split E/D with PETS/PEPSI - D/E twilight to 21:30 and 04:15-twilight

09 Feb 2021 UT - split E/D with PETS/PEPSI - E/D time twilight-20:30 and 03:00-twlight

04 Feb 2021 UT - split E/D with PETS/PEPSI - E/D until 21hr, then E/D 03:30 hr through twilight -did not open - verified elevation limit auto-configures to 30 when moving to LBCs

03 Feb 2021 UT - E/D Night - first half PETS/PEPSI - 2nd half LBTO D/E (handover 23:45) - briefly opened for PEPSI, no on-sky for D/E - took N30 LUCI-1 cals

Feb_Mar_2021_ED_Plan

2020B Technical Observing Logs

02 Jan 2021 UT - D/E Night - Observations conducted on-sky for Science D-time, Eng time executed

31 Dec 2020 UT - LBTI 1st half, E/D 2nd half - LBTI used 1st half +2 hours (bad seeing), handover to LBTO for ESM and AO eng tasks on-sky - seeing improved at end of night

30 Dec 2020 UT - LBTI 1st half, E/D 2nd half - did not open - snow - ice - Closed Dome ESM work

29 Dec 2020 UT - LBTI 1st half, E/D 2nd half - did not open - snow - ice -closed dome ESM work

28 Dec 2020 UT - E/D 2nd half - LBTI science 1st half - opened briefly first half then closed - closed dome work 2nd half

27 Dec 2020 UT - MODS AGw checked and SX/DX M3 alignments, LUCI1+PEPSI-DX transit

Dec2020_Jan2021_ED_Plan

03 Dec 2020 UT - E/D first half - PEPSI science 2nd half

02 Dec 2020 UT - E/D Night - first few hours PEPSI Pol checkout, then AO Science - late opening due to PEPSI-Pol install, seeing 2-3" highly variable - remainder of night used for PEPSI-Pol test observing

01 Dec 2020 UT - E/D Night - AO first half - GRB imaging program fairly successful, some issues with ESM longslit observations - re-rotator, open loops on offsets; power bump after midnight knocked out many systems, 2nd half LBC TMS testing in active mode

30 Nov 2020 UT - E/D Night 2nd half (1st half PEPSI) - 2nd half issue with LBC swing-arm, switched to ESM testing

Nov/Dec2020 E/D Plan

01 Nov 2020 UT - E/D Night (1st half) - split with LBTI (2nd half) - closed for time, when open too cloudy to do anything

31 Oct 2020 UT - E/D Night (1st half) - mostly PEPSI - split with LBTI (2nd half) - PEPSI starst at 17:30 local time - MODS defocus work, then LBTI

30 Oct 2020 UT - E/D Night (1st half) - split with LBTI (2nd half) - Speckle test data collected, Neptune attempt, collected some ESM data - thick clouds rolled in and out after 8:30pm local time

29 Oct 2020 UT - E/D Night (1st half) - split with LBTI (2nd half) - Speckle test (some re-rotator issues), ESM in highly variable and poor conditions (M52 bright, M52 faint failed, M34), H and K images of Uranus

28 Oct 2020 UT - E/D Night (1st half) - split with LBTI (2nd half) - No chance to open due to weather - AO testing closed dome

27 Oct 2020 UT - E/D Night (1st half) - split with LBTI (2nd half) - No chance to open due to weather - AO testing closed dome

26 Oct 2020 UT - E/D Night (1st half) - split night with LBTI (2nd half) - No chance to open due to weather - LUCI calibrations taken

October E/D Night Plan (UT Oct 26-Nov 1 2020)

02 Oct 2020 UT - Night #4 of SOUL Commissioning - 1st half LBTO / 2nd half Arcetri -

01 Oct 2020 UT - Night #3 of SOUL Commissioning - 1st half LBTO / 2nd half Arcetri - OSCO 2x; techincal observations of Pal 10, M52; power bump, Performance testing 2nd half.

30 Sep 2020 UT - Night #2 of SOUL Commissioning - 1st half LBTO / 2nd half Arcetri - tested off-axis Bin2 with previously taken on-axis Bin4 for extended objects, star cluster, ADI observations at high elevation/transit, performance tests

29 Sep 2020 UT - Night #1 of SOUL Commissioning - 1st half LBTO / 2nd half Arcetri - tested on-axis R~18, and off-axis same target, tested ADI observations at high elevation, collected science OBs, some performance tests for Bin4

28 Sep 2020 UT - Engineering NIght (includes some D-time reimbursement) - Eng/PEPSI/MODS - LBC engineering work with TMS, succesfful observations of 5 PEPSI Targets, MODS imaging of GRB down to elevation 30.

06 Sep 2020 UT - D-night / Rollover of Restart Tasks

05 Sep 2020 UT - 2020 Restart Night #15 - 1st half AO/ESM checkout continued; 2nd half LBC checkout, rotator center measured for LBCs, Superfoc seq attempted but failed due to temp drop; LUCI G200 zJ/HK observation done (possible zJ spectrum shift in telluric)

04 Sep 2020 UT - 2020 Restart Night #14 - LBTI Checkout (1st half), some success, some issues; SOUL/LUCI/SX checkout, Bin 1 and Bin 4 (R=18!!), but some elongation noted, R=12.3 AORef Bin 2 with M52 to test keepling loop closed without issues for at least 30min

03 Sep 2020 UT - 2020 Restart Night #13 - No chance to open due to lack of fuel (maintenance on fuel tanks) for backup generator, closed dome tasks only

02 Sep 2020 UT - 2020 Restart Night #12 - LUCI1/LUCI2 Rot ZP and AGW Transform collect; NS guiding LUCI1/LUCI1, LUCI1/LUCI2 imaging but some issues (IT 8179); closed for bad weather after.

01 Sep 2020 UT - 2020 Restart Night #11 - LUCI1 detector rotator center, Rotator ZP for both LUCIs (PCS updated); poor weather for transform - closed after a few hours; Closed Dome Activities: LUCI Flexure and wavelength

31 Aug 2020 UT - 2020 Restart Night #10 - No chance to open - Closed Dome Activities: LUCI-2 wavecal; LBTI Checkout; LUCI-2 cooling

30 Aug 2020 UT - 2020 Restart Night #9 - No chance to open - Closed Dome Activities: SX AO testing, LUCI2 wavelength calibrations, MODS and LBC testing

29 Aug 2020 UT - 2020 Restart Night #8 - No chance to open due to weather - Closed Dome rollover from Daytime: LUCI-1 functional checkout

28 Aug 2020 UT - 2020 Restart Night #7 - Completed AGW1 alignment, hotspot center check; Completed SX M3 alignment; LBC NS imaging; MODS TOO - MODS1B offline, LBCB offline

27 Aug 2020 UT - 2020 Restart Night #6 - LUCI-2: detector rotator center, Rotator ZP (AGW2), Imaging (Sidereal and NS); LBC-R: NS imaging, Rotational Trailing, morning skyflats

26 Aug 2020 UT - 2020 Restart Night #5 - Succssful night on-sky, completed all high priority MODS work, MODS is released for science, completed a number of LBC-R checkouts

25 Aug 2020 UT - 2020 Restart Night #4 - Did not open - some LUCI-2 functional checks, troublehsooting LBC image transfer issue, RFBG issue

24 Aug 2020 UT - 2020 Restart Night #3 - opened for about 2 hours, completed Detector Rotator Center, AGW ZP Rot Transform, and most of a spectrophotomtric std in all modes

23 Aug 2020 UT - 2020 Restart Night #2 - Did not open - ran LBC tests - LBC-Blue pulled ffrom service - MODS focus checks

22 Aug 2020 UT - 2020 Restart Night #1 - No chance to open - MODS-Bino Cals (Dual Grating and Imaging)

2020 Restart Plan - Planning Page for Aug/Sep restart

2020A Technical Observing Logs

03 Jun 2020 UT - C19 Restart Night 2 --- MODS spectrophot, LS and NS imaging obs, LBC guided & dithered obs, both sidereal & non-sidereal. LBC would not collimate, large temperature difference

02 Jun 2020 UT - C19 Restart Night 1

COVID-19 20A Restart Plan

10 Mar 2020 UT - SOUL-C Night 2

09 Mar 2020 UT - SOUL-C Night 1 --- closed dome work on SW for ESM, but primarily NCPA meas. Opened for ~2 hrs at 4am and obs M3.

08 Mar 2020 UT - Engineering Night

03 Mar 2020 UT - Unexpected E/D night - Closed for weather - Closed dome activities only (LUCI-1 and SOUL-SX)

10 Feb 2020 UT - Last night of AO Commissioning - Rigid installed on DX - checkout highest priority, SX SOUL/IRTC work

09 Feb 2020 UT - AO Commissioning night - some work completed on DX SOUL (first light achieved), DX problems later in the night - DX AdSec planned for removal next morning

08 Feb 2020 UT - Nominally an AO SOUL-COM night, DX AdSec out of use, SX IRTC2 SOUL use, D-time MODS-1 only backup - accomplished SX SOUL WFS/IRTC testing

07 Feb 2020 UT - D/E full night - IRTC2 used on-sky, some issues with: DX Adsec, PEPSI, LBCs

06 Feb 2020 UT - D/E full night - AGW2/LUCI-2 checkout completed, tested 2020B TCS, variable seeing - 2 poor weather programs attempted (1 OK)

05 Feb 2020 UT - First half D/E time-- Closed atll night - no tasks completed

04 Feb 2020 UT - First half D/E time -- Closed atll night - no tasks completed

02 Feb 2020 UT - technical time at start for AGW2/LUCI2 checks

Plan for 04-07 Feb 2020 UT - E/D nights

30 Jan 2020 UT - AGW-2 Checkout (roll over if unable to do) - procedure and script info

27 Jan 2020 UT - 1/2 Eng night (First half) - AO/ESM on-sky checks highest priority

09 Jan 2020 UT - Eng night 2nd half (split with LBTI) - confirmatory (or replacement) AGW transform data - then D-targets
08 Jan 2020 UT - Eng night 2nd half (split with LBTI) - AGW3@RFBG checkout (SH cutout issue, Rot ZP, AGW transform collection)
07 Jan 2020 UT - TSIP night 2nd half (split with LBTI) - JMH working on AGW3@RFBG checkout in parallel to LUCI/LUCI science

01 Jan 2020 UT - technical time for AGW3@RFBG checkout part trois

2019 Technical Observing Logs

31 Dec 2019 UT - technical time for AGW3@RFBG checkout redux - aligned SH spots, aligned rotator, diagnosing other issues, did not do transform or LUCI ROT zp
30 Dec 2019 UT - technical time for AGW3@RFBG checkout - did not open due to large sheets of ice dangling

Plan for 06 Dec 2019 UT - Non-ARGOS Engineering Plan for UT 06-Dec-2019 through 16-Dec-2019

17 Dec 2019 UT - TSIP/LUCI first half - ARGOS or Eng contingency 2nd half
16 Dec 2019 UT - Eng Soul in first half - ARGOS or Eng contingency 2nd half
15 Dec 2019 UT - Eng Soul in first half - ARGOS or Eng contingency 2nd half - Never opened
14 Dec 2019 UT - Eng SOUL first half - LBTI 5 hours - D-time 4am - twlight - brief opening - only LBTI obtained some data
13 Dec 2019 UT - Eng SOUL first half - ARGOS 2nd half (Eng/D Contingency) - SOUL work done, short attempt at ARGOS after high winds
12 Dec 2019 UT - Eng - SOUL first half - ARGOS 2nd half (contingeny Eng if ARGOS cannot be used)
11 Dec 2019 UT - Eng/ESM Testing (ESM is highest priority) - ESM tested (after slow start), tested single stars, open star cluster, and interacting object before clouds hit. Reverted to extreme poor weather, then backup programs before closing.

09 Dec 2019 UT - Closed Dome: ARGOS tests (Vibrations seen)
08 Dec 2019 UT - Closed Dome: tested parameter table for SX SOUL ESM. (Vibrations seen)
07 Dec 2019 UT - SHARK-VIS 1st half/E (LUCI1RZP w/M67 mask) in lieu of ARGOS 2nd half
06 Dec 2019 UT - Engineering Night, with some D-time in the middle. LUCI Rotator Zeropoints, AGW4 issues

03 Dec 2019 UT - Half Engineering Night (shared with PEPSI - LBTB time critical Local 01:35-06:35 = 1st half of night / Eng 2nd half)

20 Nov 2019 UT - Engineering Night - half night shared with LBTB (lost to weather)

18 Nov 2019 UT - Engineering Night - AGW4/LUCI-2 checkout and alignment, also MODS and PEPSIPFU

09 Nov 2019 UT - Contingency Eng Night (in case ARGOS cannot be used)
08 Nov 2019 UT - Contingency Eng Night (in case ARGOS cannot be used)
07 Nov 2019 UT - Contingency Eng Night (in case ARGOS cannot be used - plan to cover UT 07-09 Nov 2019)

14 Oct 2019 UT -Contingency Eng Night (in case ARGOS cannot be used)

13 Sep 2019 UT - Restart Night #12 - SOUL SX checkout, AGW-1 & AGW-2 transform verification, Bino LUCI(G210) spec, Bino MODS NS guiding & dithering,blind offset check, sci-like observations. LBC guiding & dithering check. LBC sky flats.
12 Sep 2019 UT - Restart Night #11 - AGW-1 transform, LBTI restart on-sky, G200 LUCI-Bino zJ,HK NS science-like observations
11 Sep 2019 UT - Restart Night #10 - Lost night due to weather - collected more LUCI flexure data
10 Sep 2019 UT - Restart Night #9 - Lost night due to weather - completed verifiction flexure LUCI N180-G210, took new flexure data N30-mirror, MODS calibrations, test of LBCs with biases, test of MODS offsets via GUI
09 Sep 2019 UT - Restart Night #8 - Lost most of night to weather - was able to test NS guiding/offsets with LUCIs. - exposure time issue still remains an issue for GCS - IT 7643
08 Sep 2019 UT - Restart Night #7 - Lost to Weather - no closed dome testing due to leaks, telescope at horizon
07 Sep 2019 UT - Restart Night #6 - LUCI-2 AGw2 Transform Data, LUCI NS guiding and dithering (GCS problem), LBC NS guiding, LBC-Superfoc, some LBC skyflats
06 Sep 2019 UT - Restart Night #5 - LUCI-1 SX pupil wobble, LUCI-Bino dither test, Field aberrations, PEPSI PFU alignment, internal LUCI-Bino JHK flats
05 Sep 2019 UT - Restart Night #4 - LUCI on-sky checks, AGW2 (wfs)/LUCI-2 taken out of service, LUCI-1/LBCR tests, AGW1/AGW2 are aligned and centered
04 Sep 2019 UT - Restart Night #3 - Bino-MODS on sky checks, completed nearly everything
03 Sep 2019 UT - Restart Night #2 - opened and closed several times, LBCs collimated, test scripts run, but LBC-B issues; MODS-2 tested but MODS-1 AGW failure; more issues see ITs
02 Sep 2019 UT - Restart Night #1 - co-pointed LBCs, attempted skyflats and collimation, lots of issues, see ITs

2019 Restart Plan - 2019 Restart Plan

09 Jul 2019 UT - Engineering Night - SOUL testing 1st half - backup Eng tasks in case of poor conditions

19 Jun 2019 UT - Engineering Night - ARGOS/Eng tasks

27 May 2019 UT - split night with LN - first half Eng - Lost to weather (wind), testing Build 2019A closed dome
26 May 2019 UT - split night with LN - first half Eng - LUCI Priority - hotspot checked, old configuration OK, AGW-1 issues, rotator ZP attempted - masks impaired completion

23 May 2019 UT - D/E Night - rollover of eng tasks - LUCI-1/AGW-1 Checkout, MODS1/2 Checkout - opened briefly, unable to collimate or point well due to TCS2019A update - troubleshooting MODS2 (cals taken)
22 May 2019 UT - D/E Night - rollover of eng tasks - LUCI-1/AGW-1 Checkout, MODS1/2 Checkout - minimal accomplished - MODS1 cals taken, troubeshooting MODS2 - never opened
21 May 2019 UT - LBTI Night - closed - used for MODS and LUCI-1 closed dome restart checks

05 May 2019 UT - Engineering/D-night - This page also serves as a starting point for UT May 5, May 17, May 22-23 D/E-nights
04 May 2019 UT - John's notes from PEPSIPOL Active Optics in parallel with PEPSIPOL science

23 Apr 2019 UT - Split with AZ (1st half) - Eng (2nd half) - weathered out - no tasks completed

10 Apr 2019 UT - SOUL Engineering 1st half (2nd half AZ - PI Conrad)
09 Apr 2019 UT - SOUL Engineering 1st half (2nd half AZ - PI Conrad)
08 Apr 2019 UT - SOUL Engineering 1st half (2nd half AZ - PI Conrad)
07 Apr 2019 UT - Engineering Night - 1st half SOUL - 2nd half Engineering
06 Apr 2019 UT - Engineering Night - 1st half SOUL - 2nd half Engineering

19 Mar 2019 UT - Engineering Night - ARGOS 50% - AZ 50% end of night
18 Mar 2019 UT - Engineering Night - ARGOS 50% - AZ 50% end of night
17 Mar 2019 UT - No Chance to Open due to Snow Intrusion Event
16 Mar 2019 UT - Engineering Night - ARGOS 75% - TOO 25% end of night
15 Mar 2019 UT - Engineering Night - ARGOS 1st half

26 Feb 2019 UT - Engineering Night - ARGOS primary, Pointing, Dithering Secondary

18 Feb 2019 UT - 1st half E, 2nd half SOUL-C (lost to weather. LBC Swing Arm exercise. MODS calib settings).
17 Feb 2019 UT - 1st half E, 2nd half SOUL-C (lost to weather. LUCI2 N30 Ks flats and MODS focus sequences done).

23 Jan 2019 UT - PEPSIPOL observing (John's notes)
21 Jan 2019 UT - PEPSIPOL observing with eclipse (John's notes)
20 Jan 2019 UT - PEPSIPOL AGW3/4 alignment cont'd
19 Jan 2019 UT - PEPSIPOL cont'd - mostly closed-dome
18 Jan 2019 UT - PEPSIPOL AGW3/4 alignment in clouds

08 Jan 2019 UT - D-time 2nd half of night

2018 Technical Observing Logs

31 Dec 2018 UT - D-time 2nd half of night

23 Dec 2018 UT - ARGOS Engineering 2nd half of night
22 Dec 2018 UT - Engineering 2nd half of night

04 Dec 2018 UT - LBTO D-Night w/E-time at end (MODS AGW Transform)

29 Nov 2018 UT - LBTO Technical/Eng Night - ARGOS/Eng tasks
28 Nov 2018 UT - LBTO Technical/Eng Night
27 Nov 2018 UT - split night Eng 1st half, AZ 2nd half - successful both halves

21 Nov 2018 UT - split night LN 1st half, ARGOS/Eng 2nd half - no joy
20 Nov 2018 UT - split night LN 1st half, ARGOS/Eng 2nd half - troubleshooting of elongation seen in LGS - systematic diagnosis and testing of each stage in the ARGOS propagation, battled thick clouds, attempted some ESM
19 Nov 2018 UT - split night LN 1st half, ARGOS/Eng 2nd half - ARGOS on-sky showed vibrations/double star images - SL OK - troubleshooting
18 Nov 2018 UT - split night LN 1st half, ARGOS/Eng 2nd half - clouds interfered - no actual closed loop ARGOS data taken
17 Nov 2018 UT - split night LN 1st half, ARGOS/Eng 2nd half - NO ARGOS activity - ESM tests run on 1 object

02 Nov 2018 UT - split E-night with SOUL-Commissioning

Night Plan for UT Oct 28-30 and Nov 1-2

30 Oct 2018 UT D/E Work - AO checkout at start, MODS, LUCI AGW confirmation
29 Oct 2018 UT D/E Work - LMIRCAM/NOMIC D-time, continue E-work, D-time at end
28 Oct 2018 UT D/E Work - Night Lost

03 Oct 2018 UT Daytime AGw4 check

24 Sep 2018 UT D/E split night with INAF
23 Sep 2018 UT D/E split night with SOUL Commissioning
22 Sep 2018 UT D/E split night with SOUL Commissioning
21 Sep 2018 UT D/E split night with SOUL Commissioning - MODS probe centering

13 Sep 2018 UT Night 12 of Restart - AGw4 collimation, LUCI filters, LUCI-wfs focus, pointing data
12 Sep 2018 UT Night 11 of Restart - AGW4 collimation, science-like observing, pointing data
11 Sep 2018 UT Night 10 of Restart - AGw4 collimation, LUCI rotator ZP, LBCs
10 Sep 2018 UT Night 9 of Restart - MODS testing all night
09 Sep 2018 UT Night 8 of Restart - LUCI AGW Transforms
08 Sep 2018 UT Night 7 of Restart - AGW4 collimation, field aberration, DX transform, LUCI center, LBCs
07 Sep 2018 UT Night 6 of Restart - pupil wobble, field aberration, LBCs
06 Sep 2018 UT Night 5 of Restart - AGW4, AGW2-4 probe alignment, LUCI dithers, LBCs
05 Sep 2018 UT Night 4 of Restart - AGW4 collimation
04 Sep 2018 UT Night 3 of Restart - AGW4, AGW2 wake-up, LBCs
03 Sep 2018 UT Night 2 of Restart - Most of night lost due to weather
02 Sep 2018 UT Night 1 of Restart - Lost due to weather

Com201809plan Plan for September 2018

26 Jun 2018 UT borrowed time - AGw3 alignment
22 Jun 2018 UT borrowed time - AGw3 alignment

20 Jun 2018 UT Eng Night - AGw3 collimation with IDL (try 5 - confirmed parity flip)
11 Jun 2018 UT Eng Night - AGw3 alignment (try 4 w/John in Austin)
09 Jun 2018 UT Eng Night - AGw3 alignment (try 3 - lost to weather)
04 Jun 2018 UT Eng Night - AGw3 alignment (wfs camera dead)
03 Jun 2018 UT Eng Night - AGw3 alignment (wfs camera dead)

22 May 2018 UT D/Eng Night
16 May 2018 UT D/Eng Night (2nd half)
05 May 2018 UT D/Eng Night

01 April 2018 UT Eng Night

31 March 2018 UT Eng Night

31 Jan 2018 UT Eng/D night: MODS1+2 transform data collected with correct MODS2 RDG

20 January 2018 UT LBC with TMS
19 January 2018 UT LBC with TMS

2017 Technical Observing Logs

28 December 2017 UT E first 1/2: LBC TMS testing, MODS2 transform data collection
26 December 2017 UT E/D first 1/2: LUCI2 ToO and collection of DX pupil wobble data with LUT (google doc)
24 December 2017 UT E first 1/2: NS AO presets through LUCI scripts

29 November 2017 UT MODS1 & MODS2 rotator centers
30 October 2017 UT MODS1 checkout (First Half)

September 2017 Technical Observing

10 October 2017 UT PEPSIPOL

20 September 2017 UT PEPSIPOL

13 September 2017 UT E time start-up Night 12
12 September 2017 UT E time start-up Night 11
11 September 2017 UT E time start-up Night 10 - first photons in PPOL, attempted LUCI transforms, PPOL guiding
10 September 2017 UT E time start-up Night 9 - PPOL on-axis guiding, LBC flats
09 September 2017 UT E time start-up Night 8 - lost to weather
08 September 2017 UT E time start-up Night 7
07 September 2017 UT E-time Start-up Night 6
06 September 2017 UT E-time Start-up Night 5
05 September 2017 UT E-time Start-up Night 4 - DX only: LUCI2 transforms, LBC-Red flats
04 September 2017 UT E-time Start-up Night 3 - LUCIs and new pointing model
03 September 2017 UT E-time Start-up Night 2 - mostly clouds, LBC continued
02 September 2017 UT E-time Start-up Night 1 - AGw and LBC wake-up, M3 alignment, NS tests

February 2017 Technical Observing

05 February 2017 UT E-time

Com201702plan Plan for February 2017

2016 Technical Observing Logs

October 2016 Technical Observing

18 October 2016 UT E-time LUCI1 AO and MODS1&MODS2 checkout
11 October 2016 UT E-time - MODS
02 October 2016 UT E-time

September 2016 Technical Observing

13 September 2016 UT Start-up Observing - Lost to weather
12 September 2016 UT Start-up Observing - UAO Slave BCU and W1 Transform
11 September 2016 UT Start-up Observing - AGw1 align, W1 Transform, Field Aberrations
10 September 2016 UT Start-up Observing - AGw1 wake-up, TCS 2016B, UAO tests CentOS 6vs7
9 September 2016 UT Start-up Observing - Lost to weather = humidity
8 September 2016 UT Start-up Observing - Lost to Hurricane Newton
7 September 2016 UT Start-up Observing - Lost to Hurricane Newton
6 September 2016 UT Start-up Observing - DX AO and LBC
5 September 2016 UT Start-up Observing - LBC

Com201609plan Plan for September 2016

Spring 2016 Technical Observing

8 July 2016 UT LUCI1-AO comm

4 July 2016 UT LUCI1-AO comm

14 June 2016 UT LUCI-1-AO tests

14 February 2016 UT E-Time
13 February 2016 UT E-Time

24 January 2016 UT MODS2 calibrations for pipeline
23 January 2016 UT TCS tests, MODS bino demo, collection of transform data for MODS1+2 and LUCIs

2015 Technical Observing Logs

December 2015 Technical Observing

05 December 2015 UT SX Adaptive Optics Tests, LFBG Field Aberrations

November 2015 Technical Observing

17 November 2015 UT Closed dome, AGw transform tests

October 2015 Technical Observing

27 October 2015 UT AGW2 alignment - good seeing
25 October 2015 UT AGW2 alignment - bad seeing
23 October 2015 UT AGW1 alignment -

September 2015 Technical Observing

Com201509plan Plan for September 2015

25 September 2015 UT LDG field data, SX PepsiPFU M3 alignment -

14 September 2015 UT Wake-up Night 12 - opened briefly, mostly lost to weather
13 September 2015 UT Wake-up Night 11 - LDG field aberrations
12 September 2015 UT Wake-up Night 10 - dofpialow
11 September 2015 UT Wake-up Night 9 - briefly dofpia and wrs
10 September 2015 UT Wake-up Night 8 - lost to weather
09 September 2015 UT Wake-up Night 7 - lost to weather
08 September 2015 UT Wake-up Night 6 - first LBC data
07 September 2015 UT Wake-up Night 5 - shutter doors, then closed for clouds

May 2015 Commissioning

13 May 2015 UT MODS2 Commissioning
12 May 2015 UT MODS2 Commissioning
11 May 2015 UT MODS2 Commissioning
6 May 2015 UT MODS2 Calibrations

April 2015 Technical Observing

5 April 2015 UT E-Time swapped
4 April 2015 UT E-Time swapped

2 April 2015 UT E-Time
1 April 2015 UT E-Time

March 2015 Technical Observing

27 March 2015 UT D-Time

24 March 2015 UT TCS tests with NS Guiding, New Active Optics, SX M1 Crosstalk
MODS2 Closed Dome Commissioning 18 March 2015 UT
11 March 2015 UT SX Rigid M2 Alignment

January 2015 Technical Observing

17 January 2015 UT D-Time

2014 Technical Observing Logs

December 2014 Technical Observing

MODS2 Commissioning Night 3 1 January 2015 UT
MODS2 Commissioning Night 2 31 December 2014 UT
30 December 2014 UT E/D-Time: MODS2 AGw Commissioning
MODS2 Commissioning Night 0 29 December 2014 UT
29 December 2014 UT D-Time
25 December 2014 UT D-Time

24 December 2014 UT D-Time

November 2014 Technical Observing

Com201411plan Plan for November 2014

MODS2 Commissioning Report 19 November 2014

24 December 2014 UT D-Time

28 November 2014 UT D-Time/E

19 November 2014 UT MODS2 AGw, LUCI1 TTM
18 November 2014 UT 2014D
17 November 2014 UT MODS2 AGw, LUCI1 TTM

September and October 2014 Technical Observing

Com201410plan Plan for October 2014

August and September 2014 Technical Observing

17 September 2014 UT Startup Night 19 - lost to weather - rain
16 September 2014 UT Startup Night 18 - lost to weather - rain
15 September 2014 UT Startup Night 17 LUCI1 transform, rotcen, etc.
14 September 2014 UT Startup Night 16 LUCI1 transform, rotcen, etc.
13 September 2014 UT Startup Night 15 - lost to weather
12 September 2014 UT Startup Night 14 MODS1+LBCR pseudo-mono + other MODS1 and LUCI1 tests
11 September 2014 UT Startup Night 13 Open dome LUCI1 tests
10 September 2014 UT Startup Night 12 - lost to weather - flexure calibration and spectral cals (after move) for LUCI1
09 September 2014 UT Startup Night 11 - lost to weather - flexure calibration and focus offset cals for LUCI
08 September 2014 UT Startup Night 10 - lost to weather - flexure validation and spectral cals for LUCI
07 September 2014 UT Startup Night 9 - lost to weather - MODS1
06 September 2014 UT Startup Night 8 Crosstalk@LFBG, Zernike Gains@LDG
05 September 2014 UT Startup Night 7 - lost to weather - rain and clouds
04 September 2014 UT Startup Night 6 - Mostly clouds
03 September 2014 UT Startup Night 5 - MODS1 NS and AGW1 wake-up and LBC WRS
02 September 2014 UT Startup Night 4 - MODS1 and LBC-Red
01 September 2014 UT Startup Night 3 - LBCs and MODS1
31 August 2014 UT Startup Night 2 - LBC-Blue
30 August 2014 UT Startup Night 1 - LBC-Blue

Com201409plan Plan for September 2014

June 2014 Technical Observing

14 June 2014 UT TCS 2014C - D / E -time
10 June 2014 UT D / E -time

April/May 2014 Technical Observing

Com201404MODSPlan MODS2 Focal Station Commissioning Plan

5 May 2014 UT MODS2 Commissioning Night 6
4 May 2014 UT MODS2 Commissioning Night 5
3 May 2014 UT MODS2 Commissioning Night 4
2 May 2014 UT MODS2 Commissioning Night 3 - lost to dust
1 May 2014 UT MODS2 Commissioning Night 2 - lost to dust and wind
30 April 2014 UT MODS2 Commissioning Night 1 - poor seeing

25 April 2014 UT Pepsi PFU
24 April 2014 UT Pepsi PFU
23 April 2014 UT Pepsi PFU: Bad Weather: Closed
22 April 2014 UT Pepsi PFU: Check of guiding and active optics: OK

18 April 2014 UT DX Field Data

March 2014 Technical Observing

22 March 2014 UT LBC IQ and ARGOS Dichroic
20 March 2014 UT ARGOS Dichroic
19 March 2014 UT AO Commissioning
18 March 2014 UT LBC Software (and LBC IQ)

February 2014 Technical Observing

27 February 2014 UT AGw2 Vignetting Tests
07 February 2014 UT Closed Dome Tests

January 2014 Technical Observing

24 January 2014 UT Technical observing

10 January 2014 UT PFU
09 January 2014 UT PFU
08 January 2014 UT PFU
07 January 2014 UT PFU

2013 Technical Observing Logs

December 2013 Technical Observing

18 December 2013 UT LFBG Field Data for AdSec
16 December 2013 UT AGW2 Transformation
14 December 2013 UT AGW2 Transformation
12 December 2013 UT AGW2 Transformation Problems

November 2013 PepsiPFU Observing

November 2013 PepsiPFU Comissioning Observing Plan DMiller

07 December 2013 UT PFU and M3 adjustment in twilight
03 December 2013 UT PFU and M3 adjustment in twilight
30 November 2013 UT PFU Setup of Guiding and Active Optics Parameters
28 November 2013 UT PFU Scheduled First Light: Setup of Guiding and Active Optics Parameters

October 2013 Technical Observing

25 October 2013 UT Rigid Secondary binodal correction, AGw2 alignment

September 2013 Technical Observing

September 2013 Technical Observing Plan JHill
Michelle's LBC Checkout List

28 September 2013 UT Parallel Technical Observing with AGw2
24 September 2013 UT LBC testing, LUCI checkout, filler science
23 September 2013 UT TCS/LBC testing, MODS checkout, filler science
22 September 2013 UT DX AO
21 September 2013 UT AGw2 Calibration and DX AO
20 September 2013 UT DX AO Commissioning
19 September 2013 UT AGw2 rough binodal correction
18 September 2013 UT DX AO tests
17 September 2013 UT Lost to weather (symbiotic closed-dome)
16 September 2013 UT DX AO tests - Mostly lost to weather
15 September 2013 UT Lost to weather (symbiotic closed-dome)
14 September 2013 UT LBTI SX
13 September 2013 UT Lost to weather (TCS 2013C closed-dome)
12 September 2013 UT LBCs and AGW2
11 September 2013 UT MODS1
10 September 2013 UT Lost to weather (mixed instrument closed-dome)
09 September 2013 UT Lost to weather
08 September 2013 UT Lost to weather
07 September 2013 UT AGW1 transform - mostly lost to clouds
06 September 2013 UT LBCs and LUCI1
05 September 2013 UT MODS1 and LBCs
04 September 2013 UT LUCI1 and AGW1 Transformation
03 September 2013 UT LUCI1 and AGW1 Calibration and AGW2
02 September 2013 UT AGW1 @ LFBG Calibration
01 September 2013 UT LBC and LDG and LFBG Wake-up

June 2013 Technical Observing

June 2013 Technical Observing Plan DMiller

25 June 2013 UT LDG Collimation tests

21 June 2013 UT LDG Collimation tests
20 June 2013 UT LFBG Collimation tests

May 2013 Technical Observing

May 2013 Technical Observing Plan DMiller

22 May 2013 UT MODS collimation issues and LBTI
21 May 2013 UT SX AdSec Realignment and LBTI

Binodal Results: 20 May 2013 Analysis of Binodal Astigmatism Measurements

16 May 2013 UT MODS binodal measurements with Rigid Secondary
13 May 2013 UT MODS binodal measurements with Rigid Secondary

April 2013 Technical Observing

April 2013 Technical Observing Plan JHill

28 April 2013 UT MODS collimation issues, Crosstalk corrections
27 April 2013 UT Modes on AdSec, Increase primary bending modes

March 2013 Technical Observing

March 2013 Technical Observing Plan JHill

30 March 2013 UT Technical Observing: DETXY, Binocular Pointing, LBTI AO
29 March 2013 UT Technical Observing: Active Optics Calibration
28 March 2013 UT Technical Observing: TCS Tests & Active Optics Calibration
27 March 2013 UT Technical Observing:
26 March 2013 UT Technical Observing:
25 March 2013 UT Technical Observing:
24 March 2013 UT Technical Observing:

February 2013 Technical Observing

January 2013 Technical Observing

Jan. 22-25 Technical Observing Plan JHill

05 Febuary 2013 UT Technical Observing:
04 Febuary 2013 UT Technical Observing: Lost to weather
03 Febuary 2013 UT Technical Observing: Open-loop tracking, verified new pointing model, clouds
02 Febuary 2013 UT Technical Observing: AGW1 w/2013A, AGW1 alignment, LFBG field data EL=40, pointing data
01 Febuary 2013 UT Technical Observing: Rotator Problems, LBC w/2013A, LBTI tests
31 January 2013 UT Technical Observing: Lost to weather
30 January 2013 UT Technical Observing: Lost to weather
25 January 2013 UT Technical Observing: Lost to weather
24 January 2013 UT Technical Observing: MODS1 Collimation, SX pointing data, Astigmatism Hysteresis
23 January 2013 UT Technical Observing: TCS 2013A-test / MODS1 Collimation
22 January 2013 UT Technical Observing: AGw#2

2012 Technical Observing Logs

December 2012 Technical Observing

Dec. 27 - Jan. 1 Technical Observing Plan JHill

01 January 2013 UT Technical Observing: Closed for Snow/Rime
31 December 2012 UT Technical Observing: Closed for Clouds and Snow
30 December 2012 UT Technical Observing - some binocular tests, but mostly lost to clouds and wind
29 December 2012 UT Technical Observing - good field aberration data for LUCI1@LFBG
28 December 2012 UT Technical Observing: Closed for Snow
27 December 2012 UT Technical Observing: Closed for Wind

November 2012 Technical Observing

Nov. 28 - Dec. 1 Technical Observing Plan DMiller

01 December 2012 UT Technical Observing
30 November 2012 UT Technical Observing
29 November 2012 UT Technical Observing. Still editing
28 November 2012 UT Technical Observing

September 2012 Technical Observing

September 7-15 Technical Observing Plan JHill

27 September 2012 UT Science Prep - Pointing Model
20 September 2012 UT AO Commissioning: night 4
19 September 2012 UT AO Commissioning: night 3
18 September 2012 UT AO Commissioning: night 2
17 September 2012 UT AO Commissioning: can't find log with either AO or Commissioning name
16 September 2012 UT AO Commissioning: repurposed to AGw#1 Startup
15 September 2012 UT Startup Night 9 - Open, MODS AGw startup, MODS and LBC startup
14 September 2012 UT Startup Night 8 - Lost to weather (humidity)
13 September 2012 UT Startup Night 7 - Lost to weather (humidity)
12 September 2012 UT Startup Night 6 - Lost to weather (rain)
11 September 2012 UT Startup Night 5 - Lost to weather (rain)
10 September 2012 UT Startup Night 4 - Mostly lost to weather (humidity and clouds)
09 September 2012 UT Startup Night 3 - Lost to weather (humidity)
08 September 2012 UT Startup Night 2 - Lost to weather (thunderstorms)
07 September 2012 UT Startup Night 1 - Lost to weather (thunderstorms)

May-June 2012 Technical Observing

May 31 through June 2 Technical Observing Plan DMiller

02 June 2012 UT Source Extractor Act Opt test and Closed AO Loop on DX
02 June 2012 (< 5:00 UT) binocular range-balancing test for collimation model
01 June 2012 UT AGw#1 Performance Data
31 May 2012 UT AdSec LUT and AGw#1 Transformation

April 2012 Technical Observing

10 April 2012 UT AGw3 and LBTI
03 April 2012 UT Laser-tracker Collimation

March 2012 Technical Observing

05 March 2012 UT AGW#3 Restart with LUCI1
04 March 2012 UT AGW#3 Restart with LUCI1

February 2012 Technical Observing

AGw#4or3 Verification Plan DMiller
AO/Pisces Plan DMiller

10 February 2012 UT AO2 commissioning
09 February 2012 UT AO2 commissioning
08 February 2012 UT AO2 commissioning - lost to weather
07 February 2012 UT AO2 commissioning
06 February 2012 UT TCS-2012B and AO/PISCES technical night (copointing offsets)
05 February 2012 UT TCS-2012B and AO/PISCES technical night
04 February 2012 UT AGw#3 Verification at LFBG
03 February 2012 UT AGw#3 Verification at LFBG: Hit Crash Limit with AGw to no work. Andrew's Laser Tracker measurements

01 February 2012 UT MODS collimation problems

January 2012 Technical Observing

LBBG Performance Verification Results, January 2012 DMiller

14 January 2012 UT LBBG Commissioning: No data during first half of night due to power outage, LBC data during second
13 January 2012 UT LBBG Commissioning: Performance verification data, Sextractor tests, lost 4 /12 hrs due to clouds and power outage
12 January 2012 UT LBBG Commissioning: Mostly lost to clouds, some performance verification data, Pisces tests, repeatability of guide probe positioning
11 January 2012 UT NON-sidereal observing/testing; LBBG Commissioning: performance verification data collected (collimation and dice5)
10 January 2012 UT NON-sidereal observing/testing; Technical Observations: testing corastig with mods
09 January 2012 UT LBBG Commissioning: Lost to weather
08 January 2012 UT LBBG Commissioning: Lost to weather
07 January 2012 UT LBBG Commissioning: New Transformation and some performance verification data collected
06 January 2012 UT Technical Observations
05 January 2012 UT Technical Observations
04 January 2012 UT LBC Technical Observations, split w/RC science

2011 Technical Observing Logs

December 2011 Technical Observing

December 2011 LBBG Plan LBBG Focal Station Commissioning and Performance Verification Plan (December)

16 December 2011 UT LBBG Commissioning:
15 December 2011 UT LBBG Commissioning: Lost to weather
14 December 2011 UT LBBG Commissioning: Lost to weather
13 December 2011 UT LBBG Commissioning: Lost to weather
12 December 2011 UT LBBG Commissioning: Lost to weather
11 December 2011 UT LBBG Commissioning: Fixed pupil wobble. Collected full transformation set
10 December 2011 UT NON-sidereal observing/testing; LBC, MODS1 and LBBG calibrations

November 2011 Technical Observing

Com201110LBBGPlan LBBG Off-axis Commissioning Plan (November)

05 November 2011 UT LBBG Commissioning: Entirely lost to wind
04 November 2011 UT LBBG Commissioning: Refined Transformation and IRTC Center of Rotation - most lost to wind and clouds
03 November 2011 UT LBBG Commissioning: Initial Transformation and Field Aberration Data at LBBG
02 November 2011 UT LBBG Commissioning: Rotator center and on-axis active optics
01 November 2011 UT LBBG Commissioning: Initial AGW1 wakeup

October 2011 Technical Observing

29 October 2011 UT Rigid Secondary installed on SX: Collimation and Pointing (2nd try)
28 October 2011 UT Rigid Secondary installed on SX: Collimation and Pointing
19 October 2011 UT Gregorian Tech: Collimation
18 October 2011 UT LBC/Gregorian Tech: Collimation
17 October 2011 UT LBC/MODS Technical Collimation Data
11 October 2011 UT More BP13 testing, AO at RFBG, LBC Pointing Data
10 October 2011 UT TCS Tests with BP13 and LBC Collimation and CoPointing
09 October 2011 UT RFBG Checkout - actual clear weather

September 2011 Technical Observing

20 September 2011 UT LDG Startup and LBC -
19 September 2011 UT LDG Startup and LBC - actual clear weather
18 September 2011 UT LFBG Startup and LBC - actual clear weather
17 September 2011 UT Bent Gregorian Startup - 1 hour on sky
16 September 2011 UT Bent Gregorian Startup - RFBG zeropoints & LFBG coma - 4 hours on sky
rain, rain, rain, clouds, rain....
4 September 2011 UT LBC collimation and Pointing
3 September 2011 UT Lost to weather

July 2011 Technical Observing Nightly Log Files

5 July 2011 UT Tuning LBC FPIA parameters.
4 July 2011 UT LBC Thermal cycling.
3 July 2011 UT Binocular Collimation and Pointing, Night 2.
2 July 2011 UT Binocular Collimation and Pointing, Night 1.

May 2011 Technical Observing Nightly Log Files

1 June 2011 UT MODS1 Commissioning & Spring SV - Some observing, second half lost to wind
31 May 2011 UT MODS1 Commissioning & Spring SV - Observed all night, good seeing, high clouds at end
30 May 2011 UT MODS1 Commissioning & Spring SV - No Observing (high winds and humidity)
29 May 2011 UT MODS1 Commissioning & Spring SV - No Observing (LDG rotator failure and high winds)
25 May 2011 UT Bias frames for MODS1 AGW prior to LBTI night

21 May 2011 UT TCS Tests BP11 / AO Preset Testing, Range Balancing
20 May 2011 UT Closed for Humidity and Ice on Building
19 May 2011 UT Lost to Weather
18 May 2011 UT TCS Tests BP11 / LBC Collimation Model Temperature Dependence
17 May 2011 UT TCS Tests BP11 / BG range balancing (mostly lost to wind)
16 May 2011 UT TCS Tests BP11 / Second half lost to wind

April 2011 Technical Observing Nightly Log Files

18 April 2011 UT LBC Technical; M1 thermal models
17 April 2011 UT MODS1 Commissioning Run 7 Night 5 - Observed all night, some cirrus and bright moon
16 April 2011 UT MODS1 Commissioning Run 7 Night 4 - Observed all night, 0.5-0.8 arcsec seeing
15 April 2011 UT MODS1 Commissioning Run 7 Night 3 - Observing through cirrus and bright moonlight
14 April 2011 UT MODS1 Commissioning Run 7 Night 2 - No Observing: High Winds all Night
13 April 2011 UT MODS1 Commissioning Run 7 Night 1 - LDG Baffle Installation, Observed All Night

March 2011 Technical Observing Nightly Log Files

Statistics on IE/CA adjustments Kboutsia

RFBG Performance Verification Plan for March 2011 UT DMiller
RFBG Performance Verification Results, March 2011 UT DMiller

25 March 2011 UT RFBG Performance Verification
24 March 2011 UT RFBG Performance Verification

23 March 2011 UT MODS1 Commissioning Run 6 Night 7 - No Observing: ice on the enclosure
22 March 2011 UT MODS1 Commissioning Run 6 Night 6 - No Observing: clouds, wind, fog, ice, and snow
21 March 2011 UT MODS1 Commissioning Run 6 Night 5 - Observed first 2h, then shutdown by high winds
20 March 2011 UT MODS1 Commissioning Run 6 Night 4 - Observed all night, some wind and clouds, active optics work
19 March 2011 UT MODS1 Commissioning Run 6 Night 3 - Observed all night with a little downtime and bright moon
18 March 2011 UT MODS1 Commissioning Run 6 Night 2 - Observed most of the night, except 3h lost to high wind
17 March 2011 UT MODS1 Commissioning Run 6 Night 1 - Observed all night: active optics and MOS first light

16 March 2011 UT RFBG Performance Verification
15 March 2011 UT RFBG Performance Verification

09 March 2011 UT LBC Active Optics calibration
08 March 2011 UT - Lost to weather
07 March 2011 UT - Lost to weather

01 March 2011 UT LBCR IQ check and DOFPIA Tests (beginning of LBTB night!)

February 2011 Technical Observing Nightly Log Files

24 February 2011 UT - Lost to weather (humidity)
23 February 2011 UT Active Optics Check out - 2nd half night MODS/LUCI active optics
22 February 2011 UT MODS1 Commissioning Run 5 Night 9 - No Observing: heavy ice accumulation on the building
21 February 2011 UT MODS1 Commissioning Run 5 Night 8 - No Observing: high humidity, ice and snow on the building
20 February 2011 UT MODS1 Commissioning Run 5 Night 7 - No Observing: high winds, humidty, and snow/ice
19 February 2011 UT MODS1 Commissioning Run 5 Night 6 - No Observing: bad weather
18 February 2011 UT MODS1 Commissioning Run 5 Night 5 - Active Optics work at start, last 6+ hours lost to high winds
17 February 2011 UT MODS1 Commissioning Run 5 Night 4 - No Observing: high winds and humidity
16 February 2011 UT MODS1 Commissioning Run 5 Night 3 - Medium to bad seeing, mostly photometric, 4h lost to high winds
15 February 2011 UT MODS1 Commissioning Run 5 Night 2 - Good seeing and photometric, lost 1h to rotator fault
14 February 2011 UT MODS1 Commissioning Run 5 Night 1 - Lost 7 hours to clouds, worked last 2 hours of the night thru cirrus

13 February 2011 UT Range Balancing second half of night
12 February 2011 UT Range Balancing second half of night

January 2011 Technical Observing Nightly Log Files

27 January 2011 UT Range Balancing second half of night

20-26 January 2011 UT AO Commissioning: Telescope disabled, retroreflector work

14-18 January 2011 UT MODS1 Commissioning Run 4 Nights 3 thru 7 - No Observing: Telescope disabled by bogie failure
13 January 2011 UT MODS1 Commissioning Run 4 Night 2 - No Observing: night lost to enclosure bogie failure
12 January 2011 UT MODS1 Commissioning Run 4 Night 1 - No observing: night lost to enclosure bogie failure

2010 Technical Observing Logs

December 2010 Technical Observing Nightly Log Files

RFBG Performance Verification Plan for December 2010 UT DMiller

28 December 2010 UT RFBG Performance Verification: Lost due to weather
27 December 2010 UT RFBG Performance Verification: rangebal - bad seeing
26 December 2010 UT RFBG Performance Verification
24 December 2010 UT RFBG Performance Verification: Lost due to weather
23 December 2010 UT RFBG Performance Verification: Lost due to weather
22 December 2010 UT RFBG Performance Verification: Lost due to weather
21 December 2010 UT RFBG Performance Verification: Lost due to weather

20 December 2010 UT MODS1 Commissioning Run 3 Night 5 - No Observing: high humidity and wind - an official no-hitter
19 December 2010 UT MODS1 Commissioning Run 3 Night 4 - No Observing: high humidity
18 December 2010 UT MODS1 Commissioning Run 3 Night 3 - No Observing: high humidity
17 December 2010 UT MODS1 Commissioning Run 3 Night 2 - No Observing: snow
16 December 2010 UT MODS1 Commissioning Run 3 Night 1 - No Observing: clouds, high winds, and high humidity
15 December 2010 UT LBC Range Balancing check-out/TCS software test

November 2010 Technical Observing Nightly Log Files

28 November 2010 UT RFBG - AO Commissioning
27 November 2010 UT RFBG - AO Commissioning
26 November 2010 UT RFBG - AO Commissioning
25 November 2010 UT RFBG - AO Commissioning: Open two hours. Rest lost to Humidity
24 November 2010 UT RFBG - AO Commissioning: Lost to High Winds
23 November 2010 UT RFBG - AO Commissioning: Open
22 November 2010 UT RFBG - AO Commissioning: Lost to Weather

21 November 2010 UT LDG - MODS1 AGw Commissioning Night 7 - high winds, high humidity all night
20 November 2010 UT LDG - MODS1 AGw Commissioning Night 6 - cloudy all night
19 November 2010 UT LDG - MODS1 AGw Commissioning Night 5
18 November 2010 UT LDG - MODS1 AGw Commissioning Night 4 - good seeing loads of transformation, field data
17 November 2010 UT LDG - MODS1 AGw Commissioning Night 3
16 November 2010 UT LDG - MODS1 AGw Commissioning Night 2
15 November 2010 UT LDG - MODS1 AGw Commissioning Night 1
14 November 2010 UT LDG - MODS1 AGw bonus time (part of TCS BP6 testing)

7 November 2010 UT AZ- LBTI - Fringes with DX AO
6 November 2010 UT AZ- LBTI - Closed DX AO loop with 153 modes
5 November 2010 UT AZ- LBTI - Adaptive Optics preparations

October 2010 Technical Observing Nightly Log Files

29 October 2010 UT AO Commissioning
28 October 2010 UT AO Commissioning
27 October 2010 UT AO Commissioning
26 October 2010 UT Open at 1:00 am MST, BP4 closed dome tests
25 October 2010 UT Open almost all night
24 October 2010 UT Open half the night
23 October 2010 UT Closed: AO Retroreflector work and Astigmatism measurements
22 October 2010 UT Closed: AO Retroreflector work and BP4 tests
21 October 2010 UT Closed: AO Retroreflector work
20 October 2010 UT Closed: AO Retroreflector work

19 October 2010 UT TCS BP4
18 October 2010 UT RFBG Checkout - Lost to clouds
17 October 2010 UT AZ- LBTI - Fringes all night long, even some at 5 microns
16 October 2010 UT AZ- LBTI - More fringes, partial loss to weather
15 October 2010 UT AZ- LBTI - First Fringes with LBTI in imaging mode at 12 microns (Hints for Co-pointing)

September 2010 Technical Observing Nightly Log Files

LFBG Re-Commissioning Plan for 14-16 September 2010 UT JHill
Binocular Testing JHill
LDG Re-Commissioning Plan with MODS1 for 22-27 September 2010 UT JHill

28 September 2010 UT LDG - MODS1 AGw - Night 6 - 1 hr lost to telescope, brief observations but most of the night time lost to high overcast
27 September 2010 UT LDG - MODS1 AGw - Night 5 - worked all night, but poor/variable seeing
26 September 2010 UT LDG - MODS1 AGw - Night 4 - 4 hrs lost due to Humidity - Rotation Center and LEFTZEROPOINT measured
25 September 2010 UT LDG - MODS1 AGw - Night 3 - mostly guider and WFS work
24 September 2010 UT LDG - MODS1 AGw - Night 2 - 5 hrs lost due to Humidity
23 September 2010 UT LDG - MODS1 AGw - Night 1 - Lost to weather (rain)
22 September 2010 UT RFBG - Lost to weather (rain)
21 September 2010 UT LFBG Field Alignment
20 September 2010 UT LBC
19 September 2010 UT LBC
18 September 2010 UT BP3 Tests / LBC
17 September 2010 UT BP3 Tests / LBC
16 September 2010 UT LFBG Re-commissioning - Collimation, Pointing, Field Aberration Data
15 September 2010 UT LFBG Re-commissioning - LEFTZEROPOINT and Transformation
14 September 2010 UT LFBG Re-commissioning - Checkout and Initial Alignment

Astigmatism Plots

RFBG Astigmatism Plots

June-July 2010 Technical Observing Nightly Log files

June-July_Technical_Schedule.pdf: June 27 - July 5

RFBG Performance Verification Plan for June/July 2010 UT DMiller

06 July 2010 UT TCS and RFBG
05 July 2010 UT RFBG
04 July 2010 UT TCS and RFBG
03 July 2010 UT RFBG
02 July 2010 UT LUCIFER
01 July 2010 UT LUCIFER
30 June 2010 UT RFBG
29 June 2010 UT RFBG
28 June 2010 UT RFBG
27 June 2010 UT Vibrations Tests

May 2010 Technical Observing Nightly Log files

24 May 2010 UT RFBG - Lost to weather
23 May 2010 UT RFBG - Lost to weather
22 May 2010 UT RFBG -

21 May 2010 UT LFBG M2 Collimation v.Andrew
20 May 2010 UT LFBG M2 Collimation v.Andrew
19 May 2010 UT LFBG M2 Collimation v.Andrew

April 2010 Technical Observing Nightly Log files

LDG Performance Verification Plan for April 2010 UT DMiller

30 April 2010 UT AO Commissioning at RFBG -
29 April 2010 UT AO Commissioning at RFBG - Lost to weather (wind)
28 April 2010 UT RFBG Commissioning - Poor conditions, some pointing data
27 April 2010 UT RFBG Commissioning - Problems first half, transformation data

LDG Performance Verification Results, April 2010 UT DMiller

26 April 2010 UT LDG Commissioning - SX Primary data, offset data, large offset data, AGw probe data
25 April 2010 UT LDG Commissioning - Preset, collimation, guiding, offset (dice5) and inverse field data
24 April 2010 UT LDG Commissioning - Lost to weather
23 April 2010 UT TCS Testing, LDG Commissioning - Lost to weather
22 April 2010 UT TCS Testing, LDG Commissioning - Lost to weather
21 April 2010 UT TCS Testing, LDG Commissioning - Lost to weather
20 April 2010 UT LDG Commissioning - 2 hours dice5 data
19 April 2010 UT LDG Commissioning - Clouds early and late, dice5 data in between

March 2010 Technical Observing Nightly Log files

RFBG Commissioning Plan for 20 March 2010 UT JHill
RFBG Performance Verification Plan DMiller
RFBG Commissioning Documents (including the above two)

06 April 2010 UT RFBG Commissioning - High winds, then W-unit tests in bad seeing
05 April 2010 UT RFBG Commissioning - Field and Transformation data plus M3 tests
04 April 2010 UT RFBG Commissioning - Field data and other tests
03 April 2010 UT RFBG Commissioning - Z6 table and W-first light
02 April 2010 UT RFBG Commissioning - Lost to weather (snow)
01 April 2010 UT RFBG Commissioning - Lost to weather (wind)
31 March 2010 UT RFBG Commissioning - Shell issues, field aberration data, transformation data
30 March 2010 UT RFBG Commissioning - Collimation issues, field aberration data, transformation data
29 March 2010 UT RFBG Commissioning - Collimation issues, transformation data, collimation model data
28 March 2010 UT RFBG Commissioning - Pointing model data, IRTC J+H ZP-data, collimation issues
27 March 2010 UT RFBG Commissioning - Lost to weather (wind)
26 March 2010 UT RFBG Commissioning - Field aberration data, collimation/coma/astigmatism issues
25 March 2010 UT RFBG Commissioning - Fiddling with collimation/coma, some transformation data, Zernike comparison
24 March 2010 UT RFBG Commissioning - Lost to weather (snow)
23 March 2010 UT RFBG Commissioning - Mostly lost to weather, but Z6 versus Elevation
22 March 2010 UT RFBG Commissioning - Active Optics and Initial Field Aberration Data
21 March 2010 UT RFBG Commissioning - Active Optics with IDL and Initial Transformation Data (0.5")
20 March 2010 UT RFBG Commissioning - Closed-dome testing of Build AO3 plus first photons on shell

08 March 2010 UT LBC/LUCIFER Tech obs - lost to weather
07 March 2010 UT LBC/LUCIFER Tech obs - lost to weather

BP1 Testing and LDG Observing Plan for 2-4 March 2010 UT DMiller

04 March 2010 UT BP1 Testing and LDG Commissioning - Completed AO2 tests with LFBG
03 March 2010 UT BP1 Testing and LDG Commissioning - AGW4 failed, but BP1 and AO2 testing with LBC
02 March 2010 UT BP1 Testing and LDG Commissioning - Lost to Weather (fog), closed-dome tests

January 2010 Technical Observing Nightly Log files

01 February 2010 UT LUCIFER prep for partner science runs - lost to bad weather
31 January 2010 UT LBC+LUCIFER prep - lost to bad weather

LDG Observing Plan for 28-29 January 2010 DMiller

29 January 2010 UT LDG Commissioning - Lost to snow
28 January 2010 UT LDG Commissioning - Lost to snow
27 January 2010 UT LFBG/LBC Commissioning - partly cloudy

04 January 2010 UT LUCIFER prep for partner science runs
03 January 2010 UT LUCIFER prep for partner science runs

Logs from past Commissioning Runs

Dec 2009 Commissioning Logs (LDG/LFBG/LUCIFER)
Nov 2009 Commissioning Logs (LDG/LFBG/LUCIFER)
Oct 2009 Commissioning Logs (LDG/LFBG/LUCIFER)
Sep 2009 Commissioning Logs (LFBG)
Jul 2009 Commissioning Logs (LUCIFER)
Jun 2009 Commissioning Logs (LFBG, LUCIFER)
May 2009 Commissioning Logs (LFBG, LDG)
Apr 2009 Commissioning Logs (LFBG)
Mar 2009 Commissioning Logs (LFBG)
Feb 2009 Commissioning Logs (LFBG, LUCIFER)
Jan 2009 Commissioning Logs (LFBG, LUCIFER)
Planning Notes, DThompson Dec 2008 Commissioning Logs (LFBG)
Nov 2008 Commissioning Logs (LFBG)
Oct 2008 Commissioning Logs (LFBG)
Aug-Sep 2008 Commissioning Logs (LFBG)
June 2008 Commissioning Logs (LFBG)
May 2008 Commissioning Logs (LFBG)
April 2008 Commissioning Logs (LFBG)

Plans and Summaries

GregorianCommissioning
AGW Commissioning Plan (680g120c, updated 08-Aug-2008) BG Technical Commissioning Plan - April (008s010b, updated 11-Apr-2009)
BG Technical Commissioning Plan - April/May (008s011b, updated 23-Apr-2009)

SummaryApril2008
WFSCstatusApril2008
AGW3 Commissioning Report (updated 26-May-2009)

AGW3 Optical Re-Alignment Report (updated 15-Jul-2010)
AGW4 Commissioning Report (updated 30-Mar-2010)

AGW4 Optical Re-Alignment Report (updated 15-Jul-2010)


#VibrationData

Vibration Data

Vibration 2008 Raw IRTC data cubes, some processed data and plots
Vibration 2009 Raw IRTC data cubes, some processed data and plots

LBC

LBC Information Information and Analysis of LBC Data such as EF

-- DavidThompson - 10 Apr 2008, revised by -- JohnHill - 15 Apr 2008
I Attachment Action Size Date Who Comment
Commissioning_Aug08.pdfpdf Commissioning_Aug08.pdf manage 37 K 28 Aug 2008 - 22:30 JohnHill  
June-July_Technical_Schedule.pdfpdf June-July_Technical_Schedule.pdf manage 77 K 24 Jun 2010 - 17:20 DougMiller June-July Technical Observing Schedule
Topic revision: r1181 - 27 Nov 2023, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback